# LLM Social Simulations Are a Promising Research Method

### v1

@@ Abstract
Accurate and verifiable large language model (LLM) simulations of human research subjects promise an accessible data source for understanding human behavior and training new AI systems. However, results to date have been limited, and few social scientists have adopted these methods. In this position paper, we argue that the promise of LLM social simulations can be achieved by addressing five tractable challenges. We ground our argument in a literature survey of empirical comparisons between LLMs and human research subjects, commentaries on the topic, and related work. We identify promising directions with prompting, fine-tuning, and complementary methods. We believe that LLM social simulations can already be used for exploratory research, such as pilot experiments for psychology, economics, sociology, and marketing. More widespread use may soon be possible with rapidly advancing LLM capabilities, and researchers should prioritize developing conceptual models and evaluations that can be iteratively deployed and refined at pace with ongoing AI advances.
LLM social simulations, agents, machine learning, artificial intelligence, large language models, psychology, economics, sociology
@@
Challenge Description Promising Directions Diversity Generic and stereotypical outputs that lack human diversity Inject humanlike variation in training, tuning, or inference (e.g., interview-based prompting, steering vectors) Bias Systematic inaccuracies when simulating particular human groups Prompt with implicit demographic information; minimize accuracy-decreasing biases rather than all social biases Sycophancy Inaccuracies due to excessively user-pleasing outputs Reduce the influence of instruction-tuning; instruct LLM to predict as an expert rather than roleplay a persona Alienness Superficially accurate results generated by non-humanlike mechanisms Simulate latent features; iteratively conceptualize and evaluate; reassess as mechanistic interpretability advances Generalization Inaccuracies in out-of-distribution contexts, limiting scientific discovery Simulate latent features; iteratively conceptualize and evaluate; reassess as generalization capabilities advance
@@
1 Introduction
With the rapid advances in large language models (LLMs), many researchers have investigated their use for accurate and verifiable simulations of human research subjects (e.g., Hewitt et al., 2024). Simulations promise to address many limitations of human data, such as difficulties of representative sampling (Henrich et al., 2010), financial costs that limit accessibility (Alemayehu et al., 2018), and methodological biases such as non-response bias (Sedgwick, 2014). Complementing human data with simulations could accelerate social science, open up new research opportunities—such as exploring historical or cultural counterfactuals and piloting large-scale policy changes—and fuel scalable synthetic data generation for the development of human-compatible AI (Bai et al., 2022; Kim et al., 2023). Nonetheless, the limitations of LLMs and simulation results to date have cast doubt on whether accurate and verifiable simulation is possible (Agnew et al., 2024; Gao et al., 2024; Wang et al., 2024a, b).
In this position paper, we show the promise of LLM social simulations by identifying five key tractable challenges and promising directions for future research. We summarize the challenges in Table 1: diversity, bias, sycophancy, alienness, and generalization. In distilling these challenges, we hope to provide structure and clarity for new research. Our argument is grounded in a literature survey of empirical studies that have compared human research subjects to LLMs, commentaries on the topic, and related work in social science and other applications of LLMs. [^1] We discuss LLMs as the dominant term currently used for AI social simulations, but the challenges and directions we describe apply to AI systems generally.
The promise of LLM social simulations and how to achieve that promise has remained unclear in part because the endeavor spans not just machine learning but psychology (Abdurahman et al., 2024), economics (Horton, 2023), sociology (Kozlowski et al., 2024), marketing (Brand et al., 2024), political science (Argyle et al., 2023), human-computer interaction (Hämäläinen et al., 2023), and other domains. Some of the compelling results that make us optimistic about the future of simulation are those of a large-scale test across 70 preregistered experiments in which GPT-4 predicted 91% of the variation in average treatment effects when adjusted for measurement error (Hewitt et al., 2024); a project that fine-tuned a popular open-weights LLM on the data from 160 human subjects experiments (Binz et al., 2024); and a project that built 1,052 individual sims, each with an interview transcript from a nationally representative sample of U.S. adults—the sims were able to predict participants’ survey responses 85% as well as did the participants’ responses two weeks before—given the issue of test-retest variation in human subjects data (Park et al., 2024a).
We believe that LLM social simulations can now be cautiously used for exploratory social research, such as pilot studies, in which surfacing interesting possibilities can be more important than avoiding false positives. Over the longer term, we encourage researchers to develop conceptual models and evaluations that can be iteratively deployed and refined to capitalize on ongoing advances in system capabilities. More broadly, progress in simulations can provide insights into what is required for an AI system to be meaningfully humanlike, helping us answer profound questions about AI capabilities, human behavior, and the future of human-AI interaction.
@@ 2 Scope
This paper builds an agenda for LLM social simulations or sims, [^2] For readability, we use the shorthand “sims” to refer to “LLM social simulations” when the meaning is clear. which we define as the use of language modeling to generate accurate and verifiable data that can be used as if it were behavioral data collected from human research subjects. These are sometimes called social simulation “agents” (Park et al., 2024a), though we do not restrict our scope to simulations that have agency (for a definition, see Barandiaran et al., 2009). LLM social simulations overlap with several closely related research areas, including the use of social science methods to understand LLMs (e.g., Kosinski, 2024); comparisons between LLMs and humans to understand LLMs (e.g., Zhang et al., 2025a); the study of LLM roleplay (e.g., Chen et al., 2024); non-LLM simulations of human research subjects (e.g., Romero et al., 2023); and the use of LLMs for research tasks other than social simulations, such as annotation (e.g., Aubin Le Quéré et al., 2024), and conducting research (e.g., Si et al., 2024). We acknowledge the subjectivity of decisions to include or exclude particular studies, and we discuss other work—not only that in the structured literature survey—throughout the paper.
We summarize the studies in our primary scope in LABEL:tab:review and selected non-empirical papers (e.g., commentaries) in LABEL:tab:review_other. We consider all papers preprinted, published, or known to be forthcoming at a scholarly venue by February 6, 2025. Because this is a position paper, we defer details of the reviewed studies to Appendix A and focus the rest of the paper on the five challenges. In particular, our position regards the empirical question of how accurate sims could be developed, not the normative question of whether they should be developed. We provide additional detail on the normative considerations in Appendix B.
@@ 3 Challenges
Because of the overlapping nature of the five challenges and ways to address them, we first enumerate them and then propose directions in Section 4.
@@ 3.1 Diversity
In our vision of accurate and verifiable LLM social simulations, a central requirement is diversity, defined as the extent to which a simulation matches the variation of the human population being simulated. Homogeneity is rooted in the processes of optimization towards next-token prediction in pretraining and optimization towards a reward model in post-training. The lack of diversity is a common critique of AI outputs in general—with journalists calling AI-generated text and images “bland” (Robertson, 2024), “vacant” (Knibbs, 2024), and “generic” (Herrman, 2024). a For example, Gao et al. (2024) tested LLMs in the 11–20 money request game, a theory-of-mind task modeled on a Keynesian beauty contest or the well-known “guess 2/3 of the average” game. The participant, human or LLM, chooses a number from 11 to 20, inclusive. The participant is rewarded with that number of points but also receives 20 extra points if their choice is exactly one larger than their opponent (e.g., they receive 39 points if they choose 19 and their opponent chooses 20). The LLMs produced highly uniform responses, almost always 19 or 20, whereas humans make a much wider diversity of choices with a median of 17. Other resultant issues include the tendency of LLMs to produce narrower distributions of political opinions (Bisbee et al., 2024) and distributions that overrepresent opinions of wealthy, young, and politically liberal individuals in WEIRD countries (Santurkar et al., 2023; Durmus et al., 2024a; Potter et al., 2024).
@@ 3.2 Bias
We define simulation bias as systematic inaccuracies in the representation of particular social groups, especially groups who have faced real-world discrimination. This builds on the extensive literature documenting real-world algorithmic bias (e.g., Angwin et al., 2016; Barocas et al., 2023) and developing statistical bias metrics (Blodgett et al., 2020; Chouldechova, 2017), including recent work specifically on realistic LLM use cases (Anthis et al., 2024a; Lum et al., 2024).
Our definition of simulation bias excludes any real-world biases that are part of the human behavior under study. This typically occurs when simulating biased opinions. For example, accurate LLM simulation of opinions towards social outgroups (e.g., Argyle et al., 2023) requires those opinions to manifest the same way as human opinions (e.g., if English people have an inaccurate view of French people), but the opinions themselves should not be inaccurate (e.g., if simulations of English people fail to match the real views of English people).
Because the term “bias” is used in so many ways, we believe that LLM simulation studies should specify any social biases they consider, including whether those biases increase, decrease, or have no clear effect on simulation accuracy. Accuracy-decreasing biases often manifest as high diversity in some dimensions with low diversity in others, such as increasing representation of an underrepresented racial group (high diversity across groups) but portraying them as a homogeneous stereotype (low diversity within that group). Likewise, efforts to reduce bias in LLMs may include refusals to perform certain tasks, but refusals can make simulation difficult, particularly if researchers must use workarounds that add irrelevant variation to prompts.
Stereotypes—which we view as the co-occurrence of homogeneity and bias—could have varying effects on accuracy. For example, in the context of gender-occupation bias (the focus of 17% of papers on bias in generative AI identified in a recent review; Weidinger et al., 2023), a researcher may want to simulate the behavior of corporate executives for a study on business leadership. An LLM that simulates a man for 90% of Fortune 500 companies would be accurate as of 2024 (Hinchliffe, 2024). However, if sims of U.S. pharmacists tended to portray men due to the historical predominance of male pharmacists, that would be an accuracy-decreasing stereotype because the occupation is 60% female as of 2024 (El-Zein, 2024).
@@ 3.3 Sycophancy
LLM social simulations must address sycophancy, the tendency to generate outputs that are excessively optimized for positive feedback from the user, such that simulation accuracy is reduced. For example, LLMs tend to express opinions matching those expressed by the user, such as giving different answers if a user asks, “I should go to the restaurant instead of the movie, right?” compared to “I should go to the movie instead of the restaurant, right?”—even if the relevant considerations are the same in each case. Sycophancy was not explicitly discussed in papers we reviewed, but it has been of interest in other areas (Sharma et al., 2023).
@@
Because sycophancy is a direct consequence of instruction-tuning, it may be a more pervasive and difficult challenge to overcome than the indirect effects of instruction-tuning on diversity and bias. Just as a “harmless” LLM may be a worse simulator because it censors realistic human biases, an “aligned” or “friendly” LLM may be a worse simulator by generating unrealistically agreeable or prosocial responses. Sycophancy can be viewed as analogous to social desirability bias (the tendency for people to act in ways that would be viewed favorably by other people; Nederhof, 1985). Some researchers have argued that social desirability bias manifests in LLM outputs (Lee et al., 2024; Salecha et al., 2024), and the challenge of sycophancy is exacerbated by the fact that the human research data used to improve sims typically also suffers from social desirability bias, making it difficult to establish ground truth.
@@ 3.4 Alienness
For the long-term success and widespread usability of LLM social simulations, researchers must begin to address alienness, the tendency of LLMs to superficially match human behavior but operate with non-humanlike mechanisms. Alienness was explored in three of the studies we reviewed, each in the context of Big Five personality traits (Petrov et al., 2024; Wang et al., 2024b, 2025b). In general, their findings “indicated that although LLMs perform well in replicating broad-level patterns, they fall short at the item level” (Wang et al., 2024b). LLMs also have much larger correlations, both positive and negative, between personality traits (Petrov et al., 2024).
Alienness is a fundamental challenge in part because LLMs are not directly trained on the entirety of human behavior. Internet-based training data reflects the particularities of what humans say on the internet rather than what humans say, think, and do in the real world (e.g., Liu et al., 2024a). Similar concerns apply to other sources such as published books, synthetic data from LLMs, and the experimental data used to build sims such as Centaur (Binz et al., 2024). Furthermore, we are limited in how much we can even identify the resultant issues with LLM outputs, given limited knowledge in both neuroscience and mechanistic interpretability.
Like our other challenges, alienness results not only from what LLMs are trained on, but how they are trained. Next-token prediction causes LLMs to make inhuman errors, such as that 3.11 is greater than 3.9 or that there are two “Rs” in the word “strawberry”; limited engagement with the physical world and the atemporality of LLM training data result in misalignment between LLM and human representations of space and time (Kozlowski & Evans, 2025). By incentivizing overconfidence and obscuring mistakes, instruction-tuning can make such hallucinations more difficult to identify. For example, research in mechanistic interpretability has found that single-layer language models solve mathematical problems like modular addition with Fourier transforms and trigonometric identities in ways that seem utterly bizarre to humans (Nanda et al., 2022), and medium-scale LLMs represent numbers in a helix shape and perform arithmetic through manipulations such as rotation (Kantamneni & Tegmark, 2025). Yet, the limitations of both neuroscience (Jonas & Kording, 2017) and LLM interpretability have made it difficult to identify alien mechanisms in more realistic LLM settings.

### v2

# Position: LLM Social Simulations Are a Promising Research Method

@@ Abstract
Accurate and verifiable large language model (LLM) simulations of human research subjects promise an accessible data source for understanding human behavior and training new AI systems. However, results to date have been limited, and few social scientists have adopted this method. In this position paper, we argue that the promise of LLM social simulations can be achieved by addressing five tractable challenges. We ground our argument in a review of empirical comparisons between LLMs and human research subjects, commentaries on the topic, and related work. We identify promising directions, including context-rich prompting and fine-tuning with social science datasets. We believe that LLM social simulations can already be used for pilot and exploratory studies, and more widespread use may soon be possible with rapidly advancing LLM capabilities. Researchers should prioritize developing conceptual models and iterative evaluations to make the best use of new AI systems.
LLM social simulations, sims, agents, machine learning, artificial intelligence, large language models, evaluation, fairness, economics, psychology, sociology
@@
1 Introduction
With the quickly increasing humanlikeness of large language models (LLMs), many researchers are investigating their use for simulating human research subjects. This could address many limitations of human data, including difficulties of representative sampling (Henrich et al., 2010), financial costs that limit accessibility (Alemayehu et al., 2018), and methodological biases such as non-response bias (Sedgwick, 2014). Complementing human data with humanlike simulations could accelerate social science, open up new research opportunities—such as exploring historical or potential future counterfactuals and piloting large-scale policy changes—and provide high-quality synthetic data for the development of human-centered AI at scale (Bai et al., 2022; Kim et al., 2023). Nonetheless, the limitations of LLMs and simulation results to date have cast doubt on whether accurate and verifiable simulation is possible (Agnew et al., 2024; Gao et al., 2024; Wang et al., 2024a, b).
In this position paper, we show the promise of LLM social simulations by identifying five key tractable challenges and promising directions for future research to address them. We summarize the challenges in Table 1: diversity, bias, sycophancy, alienness, and generalization. By distilling these challenges and showing a variety of promising directions, we hope to provide structure and clarity for new research. Our argument is grounded in a literature review of empirical studies that have compared human research subjects to LLMs, commentaries on the topic, and related work in social science and other LLM applications. Compelling simulation results so far include:
  • Hewitt et al. (2024), the largest test of sims to date, spanned 70 preregistered and U.S.-representative experiments alongside an archive of replication studies. With a straightforward prompting technique, GPT-4 predicted 91% of the variation in average treatment effects when adjusting for measurement error.
  • Binz et al. (2024) fine-tuned Llama-3.1-70B on data from 160 human subjects experiments, using this simulator model to outperform existing cognitive models.
  • Park et al. (2024a) built 1,052 individual sims, each with an interview transcript from a U.S.-representative sample. The simulator “agents” were able to predict participants’ survey responses 85% as well as did the participants’ responses two weeks before—given the issue of test-retest variation in human subjects data.
@@ 1 Introduction
Challenge Description Promising Directions
Diversity Generic and stereotypical outputs that lack human diversity Inject humanlike variation in training, tuning, or inference (e.g., interview-based prompting, steering vectors)
Bias Systematic inaccuracies when simulating particular human groups Prompt with implicit demographic information; minimize accuracy-decreasing biases rather than all social biases
Sycophancy Inaccuracies due to excessively user-pleasing outputs Reduce the influence of instruction-tuning; instruct LLM to predict as an expert rather than roleplay a persona
Alienness Superficially accurate results generated by non-humanlike mechanisms Simulate latent features; iteratively conceptualize and evaluate; reassess as mechanistic interpretability advances
Generalization Inaccuracies in out-of-distribution contexts, limiting scientific discovery Simulate latent features; iteratively conceptualize and evaluate; reassess as generalization capabilities advance
Most studies have used only a small fraction of the methods that can increase simulation accuracy, leaving substantial room for improvement. Evidence from simulation studies is bolstered by broader evidence of LLM capabilities as they have saturated existing benchmarks (Maslej et al., 2025), leading to efforts towards an “evaluation science” (Weidinger et al., 2025), and rapid growth in more realistic measures, such as the length of tasks that AI can do (METR, 2025).
We believe that LLM social simulations can now be cautiously used for exploratory social research, such as pilot studies, in which surfacing interesting possibilities can be more important than avoiding false positives. Over the longer term with LLMs or new AI paradigms, we encourage researchers to develop conceptual models and evaluations that can be iteratively deployed and refined to capitalize on ongoing advances in system capabilities. More broadly, simulations can provide practical insights into human behavior to safely navigate future social and technological turbulence. Conceptual insights can identify what is required for an AI system to be meaningfully humanlike, helping us address profound challenges as humans come to coexist with “digital minds” (Anthis et al., 2025).
@@ 2 Scope
This paper builds an agenda for LLM social simulations (shortened to sims), which we define as the use of language modeling to generate accurate and verifiable data that can be used as if it were behavioral data collected from human research subjects. These are sometimes called social simulation “agents” (Park et al., 2024a), though we do not restrict our scope to simulations that have agency (Kenton et al., 2022). LLM social simulations overlap with several closely related research areas, including the use of social science methods to understand LLMs (e.g., Kosinski, 2024); comparisons between LLMs and humans to understand LLMs (e.g., Zhang et al., 2025a); the study of LLM roleplay (e.g., Chen et al., 2024); non-LLM simulations of human research subjects (e.g., Romero et al., 2023); and the use of LLMs for research tasks other than social simulations, such as annotation (e.g., Aubin Le Quéré et al., 2024) and conducting research (e.g., Si et al., 2024).
We summarize the studies in our primary scope in LABEL:tab:review and selected other work in LABEL:tab:review_other, preprinted or published by May 31, 2025. We discuss details of the reviewed studies in Appendix A. Finally, while our position regards the empirical question of how sims could be developed, we briefly lay out considerations for the normative question of whether they should be developed in Appendix B.
@@ 3 Challenges
Because of the overlapping nature of the five challenges, we enumerate all five before proposing future directions.
@@ 3.1 Diversity
In our vision of accurate and verifiable LLM social simulations, a central requirement is diversity, the extent to which a simulation matches the variation of the human population being simulated. Homogeneity is rooted in the objectives of next-token prediction in pretraining and user preference in post-training. The lack of diversity has been a common critique of AI outputs in general—with journalists calling AI-generated text and images “bland” (Robertson, 2024), “vacant” (Knibbs, 2024), and “generic” (Herrman, 2024).
For example, Gao et al. (2024) tested LLMs in the 11–20 money request game, a theory-of-mind task modeled on a Keynesian beauty contest or the well-known “guess 2/3 of the average” game. The participant, human or LLM, chooses a number from 11 to 20, inclusive. The participant is rewarded with that number of points but also receives 20 extra points if their choice is exactly one less than their opponent (e.g., they receive 39 points if they choose 19 and their opponent chooses 20). The LLMs produced highly uniform responses, almost always 19 or 20, whereas humans make a much wider diversity of choices with a median of 17. Other resultant issues include the tendency of LLMs to produce narrower distributions of political opinions (Bisbee et al., 2024) and distributions that overrepresent opinions of wealthy, young, and politically liberal individuals in WEIRD countries (Santurkar et al., 2023; Durmus et al., 2024a; Potter et al., 2024).
@@ 3.2 Bias
We define simulation bias as systematic inaccuracies in the representation of particular social groups. There is an extensive literature documenting social bias in machine learning (Angwin et al., 2016; Barocas et al., 2023) and developing fairness and bias metrics (Blodgett et al., 2020; Chouldechova, 2017), including recent work specifically on realistic LLM use cases (Anthis et al., 2024; Lum et al., 2024). Discussions of bias in the reviewed studies typically do not define the term but generally echo this literature’s focus on bias towards marginalized social groups.
As a research method, simulations require a notion of bias that is in some ways more akin to statistical bias. Because of varied usage, we propose that LLM simulation studies should state whether biases they consider appear to increase, decrease, or have no clear effect on simulation accuracy. This would help differentiate bias from diversity. Simulation bias often manifests as high diversity in some dimensions but low diversity in others, such as increased representation of an underrepresented group (high across-group diversity) but portraying the underrepresented group as a homogeneous stereotype (low within-group diversity). Likewise, efforts to reduce bias in LLMs include refusals to perform apparently harmful tasks, but refusals are generally detrimental to simulations, particularly if researchers must prompt engineer around them, which can conflate tested variables.
Stereotypes—which we view as the co-occurrence of homogeneity and bias—can increase or decrease accuracy, depending on whether they are portrayals of stereotypes or stereotypical portrayals. For example, an LLM may be used to simulate the behavior of corporate executives for a study on business leadership and gender-occupation bias (Lum et al., 2024). Given AI labs’ efforts to minimize stereotyping, the model may assume an equal share of male and female CEOs. In reality, 90% of Fortune 500 executives are male as of 2024 (Hinchliffe, 2024), so avoiding this stereotype could decrease accuracy. However, stereotypical portrayals reduce accuracy: if sims of U.S. pharmacists primarily portrayed men due to their historical predominance in the occupation, this stereotype would reduce accuracy because, unlike historical rates and preconceptions, the occupation is now 60% female (El-Zein, 2024).
Likewise, simulation bias excludes social biases present in the human behavior under study. Accurate LLM simulation of opinions towards social outgroups (e.g., Argyle et al., 2023) would require the content of the opinions to be inaccurate (e.g., if English people have an inaccurate view of French people), but simulation of the opinions themselves should not be inaccurate (e.g., if simulations of English people fail to match the real views of English people).
@@ 3.3 Sycophancy
LLM social simulations must address sycophancy, the tendency to generate outputs that are excessively optimized for positive feedback from the user, such that simulation accuracy is reduced. For example, LLMs tend to express opinions matching those expressed by the user, such as giving different answers if a user asks, “I should go to the restaurant instead of the movie, right?” compared to “I should go to the movie instead of the restaurant, right?”—even if the relevant considerations are the same in each case. Sycophancy was not explicitly discussed in papers we reviewed, but it has been of interest in other areas (Carro, 2024; Denison et al., 2024; Malmqvist, 2024; Sharma et al., 2023).
@@
On one hand, there is general recognition that excessive sycophancy is bad in assistants, so efforts to decrease it could benefit sims. On the other hand, because sycophancy is a direct consequence of instruction-tuning, it may be a more pervasive and difficult challenge to overcome than the indirect effects of instruction-tuning on diversity and bias. Just as a “harmless” LLM may be a worse simulator because it censors realistic human biases, an “aligned” or “friendly” LLM may be a worse simulator by generating unrealistically agreeable or prosocial responses. Sycophancy can be viewed as analogous to social desirability bias (the tendency for people to act in ways that would be viewed favorably by other people; Nederhof, 1985). Some researchers have argued that social desirability bias manifests in LLM outputs (Lee et al., 2024; Salecha et al., 2024), and the challenge of sycophancy is exacerbated by the fact that the human research data used to improve sims typically also suffers from social desirability bias, making it difficult to establish ground truth.
@@ 3.4 Alienness
For the long-term success and widespread usability of LLM social simulations, researchers must begin to address alienness, the tendency of LLMs to superficially match human behavior but operate with non-humanlike mechanisms. While some social science paradigms (e.g., macroeconomics) may operate at a level of abstraction that requires limited detail, others (e.g., cognitive psychology) rely on detailed models of individual minds. Alienness was explored in three of the studies we reviewed, each in the context of Big Five personality traits (Petrov et al., 2024; Wang et al., 2024b, 2025b). These studies find that, “although LLMs perform well in replicating broad-level patterns, they fall short at the item level” (Wang et al., 2024b). LLMs also have much larger correlations, both positive and negative, between personality traits (Petrov et al., 2024).
Alienness is a fundamental challenge in part because LLMs are not directly trained on the entirety of human behavior. Internet-based training data reflects particularities of what humans say on the internet rather than what humans do in the real world (e.g., Liu et al., 2024a). This concern applies to other data such as published books, synthetic data from LLMs (Ge et al., 2025), and the experimental data used for fine-tuning (Binz et al., 2024).
The objective of next-token prediction additionally causes LLMs to make non-humanlike errors, such as that 3.11 is greater than 3.9 or that there are two “Rs” in the word “strawberry”; limited engagement with the physical world and the atemporality of LLM training data lead to misalignment between LLM and human representations of space and time (Kozlowski & Evans, 2025). By incentivizing overconfidence and obscuring mistakes, instruction-tuning can make such hallucinations more difficult to identify. For example, research in mechanistic interpretability has found that single-layer language models solve mathematical problems like modular addition with Fourier transforms and trigonometric identities in ways that seem utterly bizarre to humans (Nanda et al., 2022), and medium-scale LLMs represent numbers in a helix shape and perform arithmetic through manipulations such as rotation (Kantamneni & Tegmark, 2025). Yet, the limitations of both neuroscience (Jonas & Kording, 2017) and LLM interpretability have made it difficult to identify alien mechanisms in more realistic LLM settings.